OpenStudy (ashy98):

Identify the GCF of 9a^4b^4-27a^3b^3+18a^3b^2 A.) 9a^3b^3 B.) 9a^3b^2<---- C.) 3a^2b^3 D.) 3a^3b^2

OpenStudy (ashy98):

@Godlovesme @UnkleRhaukus @wolf1728

OpenStudy (godlovesme):

I think u r correct :) but tag other peeps to make sure

OpenStudy (ashy98):

okay thank you(:

OpenStudy (godlovesme):

Anytime : )

OpenStudy (unklerhaukus):

\[9a^4b^4-27a^3b^3+18a^3b^2\\ =9a^3b^2(ab^2)-9a^3b^2(3b)+9a^3b^2(2)\\ =9a^3b^2(ab^2-3b+2)\] yep

OpenStudy (ashy98):

Thank you !!(:

OpenStudy (08surya):

gcd is greatest common divisor that means greatest value which can divide all first talk about numbers in your prblm that is 9,18 27 9=3*3 18=2*3*3 27=3*3*3 in above you can see max coommon value in all is 3*3 that is 9 now talk about alphabets 9a^4b^4-27a^3b^3+18a^3b^2 first one=9*a*a*a*a*b*b*b*b second=9*3*a*a*a*b*b*b third=9*2*a*a*a*b*b max common term from all is 9*a^3*b^2 got it?

OpenStudy (ashy98):

yes i do thank you!!(: